# Water-rich sub-Neptunes and rocky super Earths around different Stars: Radii shaped by Volatile Partitioning, Formation, and Evolution

### v1

@@ 5 Summary and Conclusions
We investigated the masses and radii of planets resulting from coupled formation and evolution modeling under the assumption that water can be abundant on sub-Neptune and super Earth mass exoplanets. The underlying formation model relies on planetesimal accretion and includes both initially rocky and water-rich planets in the observable regime. The large water fraction on planets with masses ranging from Earth mass to tens of Earth masses is a prediction of planet formation models including orbital migration and was in a prior work found to match reasonably well the radius distribution of observed exoplanets around Solar-type stars (Burn et al., 2024). Here, we followed-up on this work and explore the mass-radius relations of these planets. Furthermore, we slightly revised the initial hydrogen and helium content and extend the simulations to lower stellar masses and discussed trends with stellar mass. In Burn et al. (2024), the water was mixed with H/He in the gaseous envelope or alternatively segregated from H/He and layered on top of the rocky core. Here, we additionally explore the effect of preferential removal of hydrogen over oxygen (fractionation) and sequestration of water into the interior of the planet.
@@
For the planets with mixed envelope composition, the location of the radius valley is increasing for stellar masses ranging from 0.1 to before remaining constant. The observed slope with orbital period is reproduced within errors. The locus of the radius valley is found to be at slightly too large radii which did not occur for the initial conditions with less H/He used in Burn et al. (2024). However, we recover the trend of longer minimal orbital periods for the sub-Neptunes compared to super Earths, which was not found in Burn et al. (2024), thanks to enforcing the presence of H/He on all planets at the end of the disk lifetime. Another iteration of modeling is required to match both aspects.
@@
The obtained mass-radius relation is in general agreement with the coupled formation-evolution modeling by Venturini et al. (2024). The differences are due to different orbital period distributions and applied limits as well as due to their pebble-accretion-informed initial conditions segregating more distinctly water-rich from rocky planets compared to the prediction of planetesimal accretion used here. As in that work, the lowest-mass volatile-rich planets are predicted at lower irradiation levels, while a cut at fixed irradiation predicts a stellar-mass independent emergence of the lowest-mass sub-Neptune.
The quantification of mass-radius characteristics suggests that a mechanism should act during formation which moderately lowers the initial water content compared to standard formation modeling (e.g. atmospheric recycling with the gaseous disk or 26Al heating and water-loss on planetesimals). Further, a reduction of the number of rocky planet building blocks is motivated. Combined, this would allow for a better match with observations by steepening the mass-radius relation of volatile-rich planets, and increasing the ratio of sub-Neptune to super Earths, which was found to be lower than found by Kepler in the nominal model. However, this conclusion relies on a heterogeneous sample of exoplanetary masses and radii and the comparison of the same model to Kepler statistics would rather indicate less mass loss (or higher metallicities). Overall, this strongly motivates the construction of a homogeneous sample of well characterized planets or occurrence rates in mass-radius-orbital period space, which is being pursued by the THIRSTEE project (Lacedelli et al., 2024).
Mass loss driven by photoevaporation is found to be mostly equally removing hydrogen and oxygen for the modeled planets because all planets with surviving volatile budgets initially contain large envelopes which require large mass loss rates. This implies, that fractionation is not necessarily occurring on sub-Neptunes and the absence of enriched (e.g. He-rich) planets cannot be used as a probe to answer the nature versus nurture question on whether evolution or formation determines the present-day exoplanet properties. Instead, occurrence or absence of fractionation in nature constrains the initial volatile budget of the planet.
@@
Sequestration of water into the interior has a large effect on the resulting mass-radius relation and should therefore be considered in future studies. The resulting mass-radius relation includes less low-density sub-Neptunes in better agreement with observations compared to the model without any interior-atmosphere coupling. However, the model with sequestration predicts a population of planets without significant primary envelopes but large water abundances. These planets would have densities similar to water-worlds which do not seem to be present in significant numbers in nature (Parviainen et al., 2024). This might indicate that less than maximum sequestration or earlier outgassing occurs. However, the model used here is likely too simplistic for a conclusive assessment and we suggest to include time-dependent solidification of the assumed magma ocean and comprehensive tracking of other chemical species and elements.

### v2

@@ 5 Summary and Conclusions
We investigated the masses and radii of planets resulting from coupled formation and evolution modeling under the assumption that water can be abundant on sub-Neptune and super Earth mass exoplanets. The underlying formation model relies on planetesimal accretion and includes both initially rocky and water-rich planets in the observable regime. The large water fraction on planets with masses ranging from Earth mass to tens of Earth masses is a prediction of planet formation models including orbital migration and was in a prior work found to match reasonably well the radius distribution of observed exoplanets around Solar-type stars (Burn2024). Here, we followed-up on this work and explore the mass-radius relations of these planets. Furthermore, we slightly revised the hydrogen and helium content after formation and extend the simulations to lower stellar masses and discussed trends with stellar mass. In Burn2024, the water was mixed with H/He in the gaseous envelope or alternatively segregated from H/He and layered on top of the rocky core. Here, we additionally explore the effect of preferential removal of hydrogen over oxygen (fractionation) and sequestration of water into the interior of the planet while improving the physical treatment of the layered scenario (i.e. by using a realistic equation of state).
@@
For the planets with mixed envelope composition, the location of the radius valley is increasing for stellar masses ranging from 0.1 to before remaining constant. The observed slope with orbital period is reproduced within errors. The locus of the radius valley is found to be at slightly too large radii and the sub-Neptune population is not numerous enough, which did not occur for the initial conditions with less H/He used in Burn2024. However, we recover the trend of longer minimal orbital periods for the sub-Neptunes compared to super Earths, which was not found in Burn2024, thanks to enforcing the presence of H/He on all planets at the end of the disk lifetime. Another iteration of modeling is required to match both aspects.
@@
The obtained mass-radius relation is in general agreement with the coupled formation-evolution modeling by Venturini2024. The differences are due to different orbital period distributions and applied limits as well as due to their pebble-accretion-informed initial conditions segregating more distinctly water-rich from rocky planets compared to the prediction of planetesimal accretion used here. As in that work, the lowest-mass volatile-rich planets are predicted at lower irradiation levels, while a cut at fixed irradiation predicts a stellar-mass independent emergence of the lowest-mass sub-Neptune.
The quantification of mass-radius characteristics suggests that at high planetary mass , a mechanism should act during formation which moderately lowers the initial water content compared to standard formation modeling. This could be a reduction of the overall budget – via atmospheric recycling with the gaseous disk or 26Al heating and water-loss on planetesimals – or by sequestration of water to the interior for which there are hints at a better match to observations at high planetary masses.
For smaller planets shaping the Kepler radius valley, assuming the photoevaporation model is accurate, metallicities intermediate between the low values of fully mixed primordial envelopes and the high values of a layered envelope or one in equilibrium with a magma ocean are required.
@@
Related to the too small locus of the radius valley, the water sequestration model predicts a population of low-mass () planets with water abundances close to unity in the gaseous envelope. These water-worlds do not seem to be present in significant numbers in nature (Parviainen2024). This might indicate that less than maximum sequestration or earlier outgassing occurs. The model used here is likely too simplistic for a conclusive assessment and we suggest to include time-dependent solidification of the assumed persistent magma ocean and comprehensive tracking of other chemical species and elements as well as their reactions (Werlen2025a) in future works. This is especially motivated by a promising match to observations at higher planetary masses.
Mass loss driven by photoevaporation on observable exoplanets is found to be mostly equally removing hydrogen and oxygen for the modeled planets because all planets with surviving volatile budgets initially contain large envelopes which require large mass loss rates. This implies, that fractionation is not necessarily occurring on sub-Neptunes and the absence of enriched (e.g. He-rich) planets cannot be used as a probe to answer the nature versus nurture question on whether evolution or formation determines the present-day exoplanet properties. Instead, occurrence or absence of fractionation in nature constrains the initial volatile budget of the planet.
The conclusions here rely on a heterogeneous sample of exoplanetary masses and radii. In addition to the discussed model improvements, we strongly encourage the construction of a homogeneous sample of well characterized planets or occurrence rates in mass-radius-orbital period space, which is, for example, being pursued by the THIRSTEE project (Lacedelli2024) and which should also be possible from the future PLATO mission and follow-up.